Find the volume and the lateral area of a frustum of a right circular cone whose radii are 4 cm and 8 cm, and the slant height is 6 cm.


1
Expert's answer
2021-02-24T06:43:28-0500

R1=4 cm, R2=8 cm, l=6 cm

Formula lateral area S=1/3"\pi"(R1+R2)l consequently S=1/3"\pi"(4+8)6=24"\pi"=75.36

Formula volume V=1/3"\pi" H (R12+R1R2+R22) for H2=l2-(R2-R1)2 consequently H="\sqrt{36-16}" =2"\sqrt{5}" and

V=1/3"\pi" 2"\sqrt{5}" (16+32+64)=224"\pi""\sqrt{5}" /3=524.25

Answer: S=24"\pi"=75.36 cm2 and V=224"\pi""\sqrt{5}" /3=524.25 cm3